Your friend tries to evaluate the idea of installing solar roof and asks for your advice. The initial costs including materials and installation are estimated to be $20,000. It will generate enough electricity for the entire household for the next 30 years. The monthly maintenance costs about $10 for the solar roof. Without the solar roof, your friend pays about $140 on electricity bill every month. Assume the annual cost of 6%. If your friend wants to have the lower monthly cost, would you suggest installing the solar roof? What is the monthly saving or additional cost if going with solar?
Answer #1

Monthly cost = 0.06 / 12 = 0.005.

PVIFA i=0.005, n = 360 = [ { 1 - ( 1 / 0.005 ) 360 } / 0.005 ] = 166.79161

Present value of the monthly cost savings = $ 130 x 166.79161 = $ 21,682.91

NPV = $ 21,682.91 - $ 20,000 = $ 1,682.91

As the NPV is greater than zero, it is suggsted that the solar toof should be installed.

Monthly saving = $ 1,682.91 / 360 = $ 4.67